Irene is a barn manager at a horse stable. She needs to calculate the correct mix of grain and hay to feed her horse. On the basis of her horse's age, weight, and workload, she determines that he needs to eat 15 lb of feed per day, with an average protein content of 8%. Hay contains 6% protein, whereas grain has a 12% content. How many pounds of hay and grain should she feed her horse each day?
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

.08*15=.06h+.12(15-h)
1.2=.06h+1.8-.12h
.12-.06h=1.8-1.2
.06h=.6
h=.6/.06
h=10 pounds of hay.
15-10=5 pounds of grain.
proof
.08*15=.06*10+.12*5
1.2=.6+.6
1.2=1.2

Answer by stanbon(48558)

calculate the correct mix of grain and hay to feed her horse. On the basis of her horse's age, weight, and workload, she determines that
he needs to eat 15 lb of feed per day, with an
average protein content of 8%.
Hay contains 6% protein, whereas
grain has a 12% content.
How many pounds of hay and grain should she feed her horse each day?
-----------------
Let weight of hay be "x" lbs; Amt of protein = 0.06x lbs
--------------------
Weight of grain is "15-x" lbs: Amt of protein = 0.12(15-x) = 1.8 -0.12x lbs
------------------
8% mixture DATA:
weight of mix 15 lbs; Amt of protein = 0.08*15 = 1.20 lbs
-------------------
EQUATION:
protein + protein = protein
0.06x + 1.8-0.12x = 1.20
-0.06x = -0.6
x= 10 lbs (amount of hay to feed the horse)
15-x=5 lbs (amount of grain to fee the horse)
================
Cheers,
Stan H.